Solve the following equation. assume that all angles are acute
Sin(3B-20`)= Cos(5B+30`)

` is my degree sign because i cant figure out how to get one

Answer by robertb(5830)

sin(3B - 20) = cos(5B + 30) = sin(90-(5B+30)) = sin(60-5B).
Then either 3B - 20 = 60- 5B,
or 3B - 20 = 180 - (60 - 5B).
The first equation gives 8B = 80, or B = 10. (A POSITIVE acute angle.)
The second equation gives 3B - 20 = 120 + 5B, or B = -70. (A NEGATIVE acute angle.)